# June 13, 2024

### v2

@@ 2.2 Disentangled Representation
$\tilde{\mathbf{x}}$ $\mathbb{R}^{d}$, reconstructed image of $\mathbf{x}$
$\mathbf{z}$$\{\mathbb{R}^{d1},…,\mathbb{R}^{dk}\}$, latent features of $\mathbf{x}$
$\mathbf{z}_{v}$$\{{\mathbb{R}^{d1},…,\mathbb{R}^{dk}}\}$, final VFs of $\mathbf{x}$
$\mathbf{z}_{a}$$\{\mathbb{R}^{d1},…,\mathbb{R}^{dk}\}$, final AFs of $\mathbf{x}$
@@
3 Proposed DifAttack++
3.1 Method Overview
Our proposed DifAttack++ aims to use disentangled VFs and AFs of images to decode AEs in score-based black-box settings, guided by the query feedback from the victim model. As illustrated in Fig. 1, DifAttack++ unfolds in two phases. In the first training stage, we train two autoencoders $\mathcal{G}$ and $\mathcal{G}^{*}$ with the same model architecture but different weights. These two models both are able to conduct feature disentanglement and image reconstruction. The training images of $\mathcal{G}$ and $\mathcal{G}^{*}$ are a set of clean image $\mathbf{x}$ and its AE $\mathbf{x}^{*}$, respectively, where $\mathbf{x}^{*}$ is generated by performing untargeted white-box attacks on surrogate models. Taking $\mathbf{x}$ as an example, $\mathcal{G}$ can disentangle the latent feature $\mathbf{z}$ of $\mathbf{x}$ into the AF $\mathbf{z}_{a}$ and VF $\mathbf{z}_{v}$ via our proposed HDF module, and then reconstructs another image $\tilde{\mathbf{x}}$ similar to $\mathbf{x}$ by decoding the fused $\mathbf{z}_{a}$ and $\mathbf{z}_{v}$. $\mathcal{G}^{*}$ also has such a same functionality for its input. In the second attack stage, we leverage the pre-trained $\mathcal{G}$ and $\mathcal{G}^{*}$ to generate an AE $\mathbf{x}^{\prime}$ of $\mathbf{x}$ to attack the black-box model $\mathcal{F}$. This $\mathbf{x}^{\prime}$ is obtained through an iterative optimization process, wherein we continuously sample a batch of perturbed images to extract AFs $\hat{\mathbf{z}}^{i}_{a}$ by updating the sampling parameters according to the query feedback, while keeping the VF $\hat{\mathbf{z}}_{v}$ consistent with that of the initial image, resulting in batches of reconstructed images $\hat{\mathbf{x}}^{i}_{f}$ to query $\mathcal{F}$, until either the query number reaches its maximum threshold $Q$ or one of the perturbed images $\hat{\mathbf{x}}^{i}_{f}$ successfully deceives $\mathcal{F}$, thereby yielding the AE $\mathbf{x}^{\prime}=\hat{\mathbf{x}}_{f}^{i}$. For clarity, a summary of notations frequently used in the following sections is provided in Table 1.
3.2 Training Stage: Train Autoencoders $\mathcal{G}$ and $\mathcal{G}^{*}$
Suppose that there is a clean image $\mathbf{x}$ in the training dataset and we have pre-trained several white-box surrogate models on this dataset, forming a set $\mathbb{C}$. Note that the availability of these white-box surrogate models has also been assumed in the existing approaches Lord et al (2022); Feng et al (2022). We adopt the white-box attack method PGD Madry et al (2018) to generate an AE $\mathbf{x^{*}}$ of $\mathbf{x}$ against a surrogate model $\mathcal{C}_{j}$ selected randomly from $\mathbb{C}$. PGD can be replaced by other white-box attack methods or a collection of them. The chosen of white-box attack methods has little impact on the final black-box attack performance, of which the comparison experiments has been presented in the supplementary file of DifAttack Liu et al (2024a). Different from our previous work DifAttack which shares one autoencoder $\mathcal{G}$ for both clean and adversarial images, in this work, we assign distinct autoencoders $\mathcal{G}$ and $\mathcal{G}^{*}$ for the clean and adversarial domain for better training efficiency and black-box attack performance (as empirically verified in the experimental section), since regular AEs usually lie off the manifold of normal images Lin et al (2020). Next, we delve into the details of the disentanglement and reconstruction tasks.
@@ 3.2 Training Stage: Train Autoencoders $\mathcal{G}$ and $\mathcal{G}^{*}$
For discovering an image’s AF to which the image’s adversarial capability is sensitive, meanwhile ensuring that alterations in this feature maintain minimal impact on the image’s visual appearance, we disentangle AFs and VFs from images’ latent feature. For realizing such a disentanglement, we exchange the AFs and VFs between a pair of clean image and its AE, expecting images reconstructed from these exchanged features to exhibit different adversarial capabilities and visual appearance according to the original images their features stem from.
HDF Module. Specifically, for implementing the feature disentanglement, as illustrated in Fig. 1(a), we design a HDF module, which inputs an image’s latent feature and disentangles this feature into VFs and AFs, finally outputing this image’s reconstructed latent feature via fusing the VFs and AFs. The HDF module consists of three kinds of convolution layers, including visual layers ($\mathbf{V}_{0}$, $\mathbf{V}_{1v}$ and $\mathbf{V}_{1a}$), adversarial layers ($\mathbf{A}_{0}$, $\mathbf{A}_{1v}$ and $\mathbf{A}_{1a}$), and fusion layers ($\mathbf{M}_{a}$, $\mathbf{M}_{v}$ and $\mathbf{M}$), which are all realized with a series of stacked $1\times 1$ convolution layers. Taking the clean image $\mathbf{x}$ as an example, HDF feeds its latent feature $\mathbf{z}=\mathcal{E}(\mathbf{x})$, where $\mathcal{E}$ denotes a Convolutional Neural Network (CNN)-based encoder, into the first disentanglement layer $\mathbf{V}_{0}$ and $\mathbf{A}_{0}$, obtaining preliminary VFs $\mathbf{z}_{v0}$ and AFs $\mathbf{z}_{a0}$ respectively as:
@@
Considering that preliminary features $\mathbf{z}_{v0}$ and $\mathbf{z}_{a0}$ obtained from disentanglement layers at one level may still be coupled, we propose to hierarchically append the second disentanglement layers $\mathbf{V}_{1v}$, $\mathbf{V}_{1a}$, $\mathbf{A}_{1v}$, and $\mathbf{A}_{1a}$ to preliminary features for achieving finer disentanglement. These four layers are used for extracting the VFs ($\mathbf{z}_{v1v}$) from the preliminary VFs ($\mathbf{z}_{v0}$) and VFs ($\mathbf{z}_{v1a}$) from the preliminary AFs ($\mathbf{z}_{a0}$), AFs ($\mathbf{z}_{a1v}$) from the preliminary VFs ($\mathbf{z}_{v0}$), and AFs ($\mathbf{z}_{a1a}$) from the preliminary AFs ($\mathbf{z}_{a0}$) as follows:
@@
Then we adopt a visual fusion layer $\mathbf{M}_{v}$ to combine these two hierarchically calibrated VFs $\mathbf{z}_{v1v}$ and $\mathbf{z}_{v1a}$, leading to the final VFs $\mathbf{z}_{v}$ of the input $\mathbf{x}$. Similarly, an adversarial fusion layer $\mathbf{M}_{a}$ is employed to obtain the final AFs $\mathbf{z}_{a}$ from $\mathbf{z}_{a1v}$ and $\mathbf{z}_{a1a}$. The generation of an image’s final VFs and AFs can be formulated as:
@@
$\mathbf{z}_{a}=\mathbf{M}_{a}\big{(}\mathbf{z}_{a1v}||\mathbf{z}_{a1a}\big{)}.$
Here, ||||| | denotes the channel-wise concatenation. Afterwards, $\mathbf{z}_{v}$ and $\mathbf{z}_{a}$ are incorporated into a reconstructed latent feature $\tilde{\mathbf{z}}$ of $\mathbf{z}$ via the fusion layer $\mathbf{M}$. Overall, we summarize the feature processing operations of HDF module as:
$\mathcal{HDF}(\mathbf{z},\mathbf{z})=\tilde{\mathbf{z}}=\mathbf{M}\big{(}\mathbf{z}_{v}||\mathbf{z}_{a}\big{)}.$ (4)
Notably, HDF does not satisfy the commutative property, i.e. $\mathcal{HDF}(\mathbf{z}^{*},\mathbf{z})\neq\mathcal{HDF}(\mathbf{z},\mathbf{z}^{*})$, where $\mathbf{z}^{*}=\mathcal{E}^{*}(\mathbf{x^{*}})$, unless $\mathbf{z}=\mathbf{z}^{*}$. To differentiate, we use $\mathcal{HDF}(\cdot)$ and $\mathcal{HDF}^{*}(\cdot)$ to represent HDF modules in networks $\mathcal{G}$ and $\mathcal{G}^{*}$, respectively. It is also easy to extend the process of $\mathcal{HDF}^{*}(\cdot)$ with an input latent feature $\mathbf{z}^{*}$ of an AE $\mathbf{x}^{*}$ as:
$\mathcal{HDF}^{*}(\mathbf{z}^{*},\mathbf{z}^{*})=\tilde{\mathbf{z}}^{*}=\mathbf{M}^{*}\big{(}\mathbf{z}_{v}^{*}||\mathbf{z}_{a}^{*}\big{)},$ (5)
where $\tilde{\mathbf{z}}^{*}$ denoted the reconstructed $\mathbf{z}^{*}$ and $\mathbf{M}^{*}$ is the fusion layer in $\mathcal{G}^{*}$. The final VFs $\mathbf{z}_{v}^{*}$ and AFs $\mathbf{z}_{a}^{*}$ of $\mathbf{x}^{*}$ are obtained through the same process as in E.q. (1)-(3), with the only difference being the substitution of modules of $\mathcal{G}$ with modules of $\mathcal{G}^{*}$.
Feature Interchange. After elucidating the HDF module, we will now introduce how to utilize HDF to conduct feature interchange for accomplishing disentanglement. Specifically, as depicted by the yellow path in Fig. 1(a), the AF of the clean sample, i.e. $\mathbf{z}_{a}$, is firstly concatenated with the VF of the adversarial image, i.e. $\mathbf{z}_{v}^{*}$. Afterward, they are inputted into the fusion layer $\mathbf{M}$ to obtain the fused feature $\mathbf{z}_{f}$. Symmetrically, as shown by the blue path, we can obtain another fused feature $\mathbf{z}_{f}^{*}$. These two steps are represented as:
@@
Here, $\mathcal{D}$ and $\mathcal{D}^{*}$ denote the CNN-based decoder in $\mathcal{G}$ and $\mathcal{G}^{*}$, respectively.
Disentanglement Loss. Subsequently, since we desire the AF to control the adversarial capability of the image, while the VF to govern its visual perception, $\mathbf{x}_{f}$ should be closer to the visual perception of the AE $\mathbf{x^{*}}$ and be non-adversarial. Conversely, $\mathbf{x}_{f}^{*}$ is expected to be closer to the visual perception of the clean image $\mathbf{x}$ and be adversarial against all surrogate models so as to improve the generalization ability of the AF on different classifiers. Hence, we aim to minimize the disentanglement loss ${L}_{dis}$ that is defined as:
@@
$\max\left\{\mathbb{I}(v)\cdot\big{(}\mathcal{C}_{j}(\mathbf{x},y)-\max\limits_{d\neq y}\mathcal{C}_{j}(\mathbf{x},d)\big{)},-k\right\}.$
The value of $v$ can be set as 0 or 1. $\mathbb{I}(0)=1$ means an untargeted attack with a ground-truth label of $y$, while $\mathbb{I}(1)=-1$ indicates a targeted attack with a target category of $y$. $\mathcal{C}_{j}(\mathbf{x},d)$ is the output score of $\mathcal{C}_{j}$ at the $d$-th class. The adjustable constant $k>0$ controls the adversarial extent.
@@
For achieving a cycle consistency Zhu et al (2017) with regard to feature disentanglement, as depicted in the reconstruction paths of Fig. 1(a), the autoencoder is expected to reconstruct an input image well even when its latent feature is processed by the HDF module. Therefore, the output of HDF modules, i.e. $\tilde{\mathbf{z}}$ and $\tilde{\mathbf{z}}^{*}$, are expected to reconstruct $\tilde{\mathbf{x}}\approx\mathbf{x}$ and $\tilde{\mathbf{x}}^{*}\approx\mathbf{x}^{*}$ through:
$\tilde{\mathbf{x}}=\mathcal{D}(\mathcal{HDF}(\mathbf{z},\mathbf{z})),$ (10)
$\tilde{\mathbf{x}}^{*}=\mathcal{D}^{*}(\mathcal{HDF}^{*}(\mathbf{z}^{*},\mathbf{z}^{*})).$
@@
Algorithm 1 The proposed DifAttack++ method.
Input: target classifier $\mathcal{F}$; autoencoder $\mathcal{G}^{*}$ pretrained in the adversarial domain; clean image $\mathbf{x}$; initial image $\hat{\mathbf{x}}$; max query budget $Q>0$; distortion budget $\epsilon$; ground-truth/target label $y$ for untargeted/targeted attacks; learning rate $\eta$; variance $\sigma$; $v$, $k$ in Eq.(LABEL:eq:algo3); sample scale $\tau$.
@@
6:        Obtaining $\hat{\mathbf{x}}_{f}^{i}$ using E.q. (LABEL:eq:init), (LABEL:eq:attackobjpre), (LABEL:eq:adv1), and (17).
@@
8:        Obtaining $\hat{\mathbf{x}}_{f}^{i}$ using E.q. (21), (LABEL:eq:open1), (LABEL:eq:adv1), and (17).

### v3

@@ 3.1 Training Stage: Train Autoencoders $\mathcal{G}$ and $\mathcal{G}^{*}$
$\mathcal{E}$ Encoder in $\mathcal{G}$
$\mathcal{D}$ Decoder in $\mathcal{G}$
@@
$\tilde{\mathbf{x}}$ $\mathbb{R}^{d}$, Reconstructed image of $\mathbf{x}$
@@
$\hat{\mathbf{x}}_{f}^{i}$ $\mathbb{R}^{d}$, $i$th reconstructed perturbed image
$\delta$ $\mathbb{R}^{d}$, Image perturbation
$\mathbf{z}$ Latent features of $\mathbf{x}$
$\tilde{\mathbf{z}}$ Reconstructed latent features of $\mathbf{x}$
$\mathbf{z}_{v}$ Final VFs of $\mathbf{x}$
$\mathbf{z}_{a}$ Final AFs of $\mathbf{x}$
$\mathbf{z}_{v0}$ Preliminary VFs of $\mathbf{x}$
$\mathbf{z}_{a0}$ Preliminary AFs of $\mathbf{x}$
$\mathbf{z}_{v1v}$ VFs of $\mathbf{z}_{v0}$
$\mathbf{z}_{v1a}$ VFs of $\mathbf{z}_{a0}$
$\mathbf{z}_{a1v}$ AFs of $\mathbf{z}_{v0}$
$\mathbf{z}_{a1a}$ AFs of $\mathbf{z}_{a0}$
$\hat{\mathbf{z}}_{a}^{i}$ Final AFs of $i$th perturbed image
$\hat{\mathbf{z}}_{v}$ Final VFs of image $\hat{\mathbf{x}}$
$\mathbf{z}_{f}$ Fused feature after feature interchange
$\mathcal{Q}$ Maximum query number
$\mathbb{T}$ Training dataset
$\mathbb{C}$ Set of white-box surrogate models
$\mathcal{C}_{j}$ $j$th white-box surrogate model in $\mathbb{C}$
$N$ Number of surrogate models in $\mathbb{C}$
$\mathcal{A}$ White-box attack method
@@
For discovering an image’s AF to which the image’s adversarial capability is sensitive, meanwhile ensuring that alterations in this feature maintain minimal impact on the image’s visual appearance, we disentangle AFs and VFs from images’ latent feature. To this end, we exchange the AFs and VFs between a pair of clean image and its AE. Then two images reconstructed from these exchanged features are expected to exhibit different adversarial capabilities and visual appearance, according to the original images their features stem from.
HDF Module. Specifically, for implementing the feature disentanglement, as illustrated in Fig. 1(a), we design an HDF module, which accepts an image’s latent features as inputs and disentangles it into VFs and AFs. Eventually, HDF also generates this image’s reconstructed latent feature via fusing the VFs and AFs. HDF module consists of three kinds of convolution layers, including visual layers ($\mathbf{V}_{0}$, $\mathbf{V}_{1v}$ and $\mathbf{V}_{1a}$), adversarial layers ($\mathbf{A}_{0}$, $\mathbf{A}_{1v}$ and $\mathbf{A}_{1a}$), and fusion layers ($\mathbf{M}_{a}$, $\mathbf{M}_{v}$ and $\mathbf{M}$), which are all implemented with a series of stacked $1\times 1$ convolution layers. Taking the clean image $\mathbf{x}$ as an example, HDF feeds its latent feature $\mathbf{z}=\mathcal{E}(\mathbf{x})$, where $\mathcal{E}$ denotes a Convolutional Neural Network (CNN)-based encoder, into the first disentanglement layer $\mathbf{V}_{0}$ and $\mathbf{A}_{0}$, obtaining preliminary VFs $\mathbf{z}_{v0}$ and AFs $\mathbf{z}_{a0}$ respectively as:
@@
Considering that preliminary features $\mathbf{z}_{v0}$ and $\mathbf{z}_{a0}$ obtained from disentanglement layers at one level may still be coupled, we propose to hierarchically append the second disentanglement layers $\mathbf{V}_{1v}$, $\mathbf{V}_{1a}$, $\mathbf{A}_{1v}$, and $\mathbf{A}_{1a}$ to preliminary features for achieving the finer disentanglement. These four layers are used for extracting the fine-grained VFs or AFs from preliminary VFs and AFs as follows:
@@
Then we adopt a visual fusion layer $\mathbf{M}_{v}$ to combine these two hierarchically calibrated VFs $\mathbf{z}_{v1v}$ and $\mathbf{z}_{v1a}$, leading to the final VFs $\mathbf{z}_{v}$ of the input $\mathbf{x}$. Similarly, an adversarial fusion layer $\mathbf{M}_{a}$ is employed to obtain the final AFs $\mathbf{z}_{a}$ from $\mathbf{z}_{a1v}$ and $\mathbf{z}_{a1a}$. Such a generation process can be formulated as:
@@
$\mathbf{z}_{a}=\mathbf{M}_{a}\big{(}\mathbf{z}_{a1v}||\mathbf{z}_{a1a}\big{)},$
where ||||| | denotes the channel-wise concatenation. Afterwards, $\mathbf{z}_{v}$ and $\mathbf{z}_{a}$ are incorporated into a reconstructed latent feature $\tilde{\mathbf{z}}$ of $\mathbf{z}$ via the fusion layer $\mathbf{M}$. Overall, we summarize the feature processing operations of the HDF module as:
$\mathcal{HDF}(\mathbf{z})=\mathbf{M}\big{(}\mathbf{z}_{v}||\mathbf{z}_{a}\big{)}=\tilde{\mathbf{z}}.$ (4)
For the clarity, we use $\mathcal{HDF}(\cdot)$ and $\mathcal{HDF}^{*}(\cdot)$ to represent HDF modules in networks $\mathcal{G}$ and $\mathcal{G}^{*}$, respectively. It is also easy to extend the process of $\mathcal{HDF}^{*}(\cdot)$ with an input latent feature $\mathbf{z}^{*}$ of an AE $\mathbf{x}^{*}$ as:
$\mathcal{HDF}^{*}(\mathbf{z}^{*})=\mathbf{M}^{*}\big{(}\mathbf{z}_{v}^{*}||\mathbf{z}_{a}^{*}\big{)}=\tilde{\mathbf{z}}^{*},$ (5)
where $\tilde{\mathbf{z}}^{*}$ denoted the reconstructed $\mathbf{z}^{*}$ and $\mathbf{M}^{*}$ is the fusion layer in $\mathcal{G}^{*}$. The final VFs $\mathbf{z}_{v}^{*}$ and AFs $\mathbf{z}_{a}^{*}$ of $\mathbf{x}^{*}$ are obtained through the same process as in Eqs. (1)-(3), with the only difference being the substitution of modules of $\mathcal{G}$ with modules of $\mathcal{G}^{*}$.
Feature Interchange in CD. After elucidating the HDF module, we now introduce how to utilize HDF to conduct the feature interchange in CD for accomplishing the disentanglement. Specifically, as depicted by the yellow path in Fig. 1(a), the AF of the clean sample (clean domain), i.e. $\mathbf{z}_{a}$, is firstly concatenated with the VF of the adversarial image (adversarial domain), i.e. $\mathbf{z}_{v}^{*}$. Afterwards, they are fed into the fusion layer $\mathbf{M}$ to obtain the fused feature $\mathbf{z}_{f}$. Symmetrically, as shown by the blue path, we can obtain another fused feature $\mathbf{z}_{f}^{*}$. These two steps can be expressed as:
@@
where $\mathcal{D}$ and $\mathcal{D}^{*}$ denote the CNN-based decoder in $\mathcal{G}$ and $\mathcal{G}^{*}$, respectively.
Disentanglement Loss. Subsequently, since we desire the AF to control the adversarial capability of the image, while the VF to govern its visual perception, $\mathbf{x}_{f}$ should be closer to the visual perception of the AE $\mathbf{x^{*}}$ and be non-adversarial. Conversely, $\mathbf{x}_{f}^{*}$ is expected to be closer to the visual perception of the clean image $\mathbf{x}$, and be adversarial against all surrogate models so as to improve the generalization ability of the AF on different classifiers. Hence, we aim to minimize the disentanglement loss ${L}_{dis}$ that is defined as:
@@
$\max\left\{\mathbb{I}(v)\cdot\big{(}\mathcal{C}_{j}(\mathbf{x},y)-\max\limits_{h\neq y}\mathcal{C}_{j}(\mathbf{x},h)\big{)},-k\right\}.$
Here, the value of $v$ can be set as 0 or 1. $\mathbb{I}(0)=1$ means an untargeted attack with a ground-truth label of $y$, while $\mathbb{I}(1)=-1$ indicates a targeted attack with a target category of $y$. Also, $\mathcal{C}_{j}(\mathbf{x},h)$ is the output score of $\mathcal{C}_{j}$ for the $h$-th class. The adjustable constant $k>0$ controls the adversarial extent.
@@
For achieving a cycle consistency Zhu et al (2017) with regard to the feature disentanglement, as depicted in the reconstruction paths of Fig. 1(a), the autoencoder is expected to reconstruct an input image well even when its latent feature is processed by the HDF module. Therefore, the output of HDF modules, i.e. $\tilde{\mathbf{z}}$ and $\tilde{\mathbf{z}}^{*}$, are expected to reconstruct two images $\tilde{\mathbf{x}}\approx\mathbf{x}$ and $\tilde{\mathbf{x}}^{*}\approx\mathbf{x}^{*}$ through:
$\tilde{\mathbf{x}}=\mathcal{D}(\mathcal{HDF}(\mathbf{z})),$ (10)
$\tilde{\mathbf{x}}^{*}=\mathcal{D}^{*}(\mathcal{HDF}^{*}(\mathbf{z}^{*})).$
@@
Algorithm 1 The proposed DifAttack++.
Input: Target classifier $\mathcal{F}$; autoencoder $\mathcal{G}^{*}$ pretrained in the adversarial domain; clean image $\mathbf{x}$; initial image $\hat{\mathbf{x}}$; max query budget $Q$; distortion budget $\epsilon$; ground-truth/target label $y$ for untargeted/targeted attacks; learning rate $\eta$; variance $\sigma$; $v$, $k$ in Eq.(LABEL:eq:algo3); sample scale $\tau$.
@@
6:        Obtaining $\hat{\mathbf{x}}_{f}^{i}$ using Eqs. (LABEL:eq:init), (LABEL:eq:attackobjpre), (LABEL:eq:adv1), and (17).
@@
8:        Obtaining $\hat{\mathbf{x}}_{f}^{i}$ using Eqs. (21), (LABEL:eq:open1), (LABEL:eq:adv1), and (17).